OpenStudy (anonymous):

Two bicycles depart from Miami Beach going in opposite directions. The first bicycle is traveling at 10 miles per hour. The second bicycle travels at 5 miles per hour. How long does it take until the bikes are 45 miles apart?

OpenStudy (dumbcow):

d = r*t total distance apart = 10t + 5t = 15t -> 15t = 45 t = 3

OpenStudy (mertsj):